In a survey of children’s favorite fruit juices, these were the result
Apple=25% Orange=14% Grape=30% Mango=31%
20 more children chose grape than chose apple.
How many children took part in the survey?

Answer by ikleyn(53765)

.

Let x be the total number of children.


Then the equation is


    0.3x - 0.25x = 20


saying that "20 more children chose grape than chose apple".


From the equation,


    0.05x = 20,

    x     = 20/0.05 = 2000/5 = 400.


ANSWER.  400 children.